9000 dollars is placed in an account with an annual interest rate of 8%. How much will be in the account after 17 years, to the nearest cent?


1
Expert's answer
2021-12-20T03:57:52-0500

"R=P(1+i)^n" , where P - initial sum, i - interest rate, n - number of payments

Since the rate is annual, then there will be 17 payments in 17 years, so n =17. i is given 0.08 and P is given 9000, so

"R=9000*(1+0.08)^{17}=9000*1.08^{17}\approx33000.16" dollars